OpenStudy (anonymous):

write the equation of the line that has a slope of a -3/4 and passes through (-2,1)

OpenStudy (anonymous):

you may use the y - y1 = (slope)(x-x1). example: slope = 1/3, and point is (0,1). so,

OpenStudy (anonymous):